A carbon compod contains 12.8% carbon,2.1%hydrogen, 85.1% bromine. The molecular weight of the compound is 187.9. calculate the molecular formula

Grade:11

1 Answers

varun
6 years ago
HELLO AASHRITH
i will explain the above problem in the following steps
STEP 1
percentage composition present in the compound is
Carbon = 12.8
Hydrogen = 2.1
Bromine = 85.1
STEP 2
Dividing with thier respective atomic weights of the elements, we get
12.8/12 =1.067
2.1/1 =2.1
85.1/80 =1.067
STEP 3
Dividing by smallest number to get simple atomic ratio,we get
1.067/1.067 =1
2.1/1.067 =2
1.067/1.067 =1
the empirical formula is CH2Br.
So,Empirical formula weight =12+(2x1)+80
=94
molecular weight is given =187.9
so
n=187.9/94
=2
So molecular formula =2xempirical formula
                                   =CH2Br x 2
                                    = C2H4Br2
so this is the answer
THANK YOU